OpenStudy (anonymous):

I really need help. 7. The function t(x) = 2x + 4 determines how many cans of tomato sauce a food truck needs to stock, where x is the number of shifts the crew is going to work in the truck. The crew uses c(t(x)) to find the amount of money to spend on tomato sauce. The function c(x) = 3x + 1. Solve for how much money must be spent when the crew is going to work 2 shifts. A. 7 B. 8 C. 18 D. 25

OpenStudy (phi):

This takes 2 steps: they tell you t(x) = 2x + 4 where x is the number of shifts crew is going to work 2 shifts first figure out t(2). can you do that ?

OpenStudy (anonymous):

t=x+2

OpenStudy (anonymous):

or t=4

OpenStudy (anonymous):

@phi

OpenStudy (phi):

almost, t(x) = 2*x + 4 to figure it out, erase the x and put in a 2

OpenStudy (phi):

in other words, you used t(x)=x+2 but that is not the equation they gave you.

OpenStudy (anonymous):

So, t(2)=2(2)+4

OpenStudy (phi):

yes, now simplify

OpenStudy (anonymous):

t=4

OpenStudy (phi):

t(2) = 2(2)+4 which means t(2) is 2*2 + 4 what is 2*2 + 4 ?

OpenStudy (anonymous):

8

OpenStudy (phi):

yes

OpenStudy (phi):

t(2) = 8 we are done with that step next step. The crew uses c(t(x)) which means they want to figure out c(8)

OpenStudy (phi):

t(2) does not mean t*2 when they write t(x)= 2x + 4 they are saying: 2x+4 is a formula, and to keep things short, we give a NAME to the formula. we call it formula "t" we put (x) next to the t: t(x) to remind us that the formula uses "x"

OpenStudy (phi):

c(x) = 3x + 1 is a different formula, which we call formula "c" we want to use this new formula with x=8

OpenStudy (anonymous):

So, for C x=8 C(x)=3x+1 C(8)=3(8)+1 C(8)=24+1 C(8)=25

OpenStudy (phi):

exactly

OpenStudy (anonymous):

Is that the answer. D. 25

OpenStudy (phi):

to review what we did The crew uses c(t(x)) to find the amount of money the c( t(x) ) is a bit complicated, but it means use formula c with x replaced by t(x) what does t(x) mean? it means use formula t. what x should we use luckily they tell us x is the # of shifts, and they want 2 shifts so we want to figure out t(2), which means use formula t with x replaced by 2 then we use that number in c

OpenStudy (phi):

yes, 25 is the answer.

OpenStudy (anonymous):

that you

OpenStudy (phi):

yw